OpenStudy (anonymous):

simplify (5x2 + 3x + 4) − (2x2 − 6x + 3). If the final answer is written in the form Ax2 + Bx + C, what is the value of B? I think B is -3 am I correct?!

OpenStudy (anonymous):

@whpalmer4 3x^2 + -3x + 7 & B is -3 am I right?! :)

OpenStudy (mathstudent55):

Can you rewrite the expression without parentheses?

OpenStudy (anonymous):

@mathstudent55 yes..

OpenStudy (mathstudent55):

What do you get? Hint: The first set of parenthese are not needed, so you can just drop them. The second set of parentheses have a negative sign to the left, so the negative has to be distributed by every term inside the parentheses.

OpenStudy (anonymous):

Wait, I already did all the work, & I just need to know if I did it right or not.. I put my answer up.. if you didn't see.. really hope this didn't sound rude lol I'm not trying to be lol @mathstudent55 :/

OpenStudy (mathstudent55):

Ok, than answer is no. That is not the right value for B.

OpenStudy (mathstudent55):

This is why I aksed to see how you simplified the expression. You need to drop the parentheses, but you need to be careful with the second set of parentheses becuase of the negative sign to its left.

OpenStudy (anonymous):

alright! I'll do what you said then! I did this problem with the vertical method & used a chart! So hold on let me redo it! give me a minute! @mathstudent55

OpenStudy (mathstudent55):

Here's the original problem, ok? (5x2 + 3x + 4) − (2x2 − 6x + 3)

OpenStudy (anonymous):

okay so I redid it & I think its either -9 or 9! .. lol @mathstudent55

OpenStudy (mathstudent55):

Since there is no number or negative sign to the left of the first set of parentheses, you can just drop the first seto of parentheses. 5x2 + 3x + 4 − (2x2 − 6x + 3) Now how do you get rid of the second set of parentheses?

OpenStudy (mathstudent55):

Your new answer is better. Now we're going to figure out if B = 9 or B = -9.

OpenStudy (anonymous):

I think you have to put a 1 outside the parentheses.. I'm leaning towards -9, but I really wanna learn how to do this.. I don't just want you to give me the answer! :) @mathstudent55

OpenStudy (mathstudent55):

5x2 + 3x + 4 − (2x2 − 6x + 3) A negative to the left of parentheses, is as if you had -1 to the left of parentheses: 5x2 + 3x + 4 − 1(2x2 − 6x + 3)

OpenStudy (mathstudent55):

Now you need to distribute the -1 by all terms inside the parentheses.

OpenStudy (mathstudent55):

The first three terms from the first set of parentheses stay the same. The negative 1 only affects what is insdide the second set of parentheses.

OpenStudy (anonymous):

okay so what do you do after that? After you distribute this is where I get confused.. lol :/@mathstudent55

OpenStudy (mathstudent55):

You need to multiply -1 * 2x^2 -1 * (-6x) -1 * 3

OpenStudy (anonymous):

& then you'd get your answer? @mathstudent55

OpenStudy (mathstudent55):

Then you combine like terms to get the answer.

OpenStudy (anonymous):

okay & it'd be -9 right?! I'm sorry if its wrong I'm really bad at math! @mathstudent55

OpenStudy (mathstudent55):

Did you do these multiplications? -1 * 2x^2 -1 * (-6x) -1 * 3

OpenStudy (anonymous):

no its positive 9!

OpenStudy (mathstudent55):

right

OpenStudy (anonymous):

alright thanks!! :)

OpenStudy (mathstudent55):

-1 * 2x^2 = -2x^2 -1 * (-6x) = 6x -1 * 3 = -3 5x2 + 3x + 4 − 1(2x2 − 6x + 3) = 5x^2 + 3x + 4 - 2x^2 + 6x - 3 = 3x^2 + 9x + 1 = Ax^2 + Bx + c B = 9

OpenStudy (mathstudent55):

wlcm